Privacy and Security Notice

FEL Upgrade Project Weekly Brief
September 24-28, 2001


We expect to receive tonight at 7:00pm a major deliverable for the FEL Upgrade Project:  the two 100 kW
klystrons for the upgrade injector.  After passing all qualifications at the vendor, CPI in Palo Alto earlier this
month, the two klystrons were shipped to us last night.

We placed our first contract for the manufacture of dipole magnets (GW-the optical chicanes), and we have
excellent results on the magnetic measurement of the wiggler.  Therefore, all major deliverables for the Phase 1
contract which ends on Sept. 30th are in:  the RF systems; the wiggler, and reviewed designs and first
components on the optical and electron beam transport systems.

Today we completed a successful 2-week run for the Joint Technology Office on laser-material damage studies.
This completes the full exposure matrix that was presented in the JTO proposal. For details see the operations

The science and technology involved in the Jefferson Lab FEL program was reviewed this week on Sept.
24-26th during the annual DOE Review of Jefferson Lab’s programs.  The FEL program received high marks for
meeting and exceeding technical goals, project planning , and the high quality science that has been demonstrated
by the FEL users over the past year.  The reviewer’s also took note of the important synergy among the lab’s
advanced accelerator R&D efforts and the lab’s Navy funded FEL upgrade program and DOE-funded linac
contract for the Spallation Neutron Source.

We completed the project monthly reports for the month of August and they will be distributed to the DOE and
ONR contract monitors on Monday.  We anticipate a smooth close of the financials for the Phase 1 contract
whose performance period ends on Sunday.  We thank the lab’s finance and procurement staff for their efforts.

The “Work for Others” contract for the Aerospace portion of the UV FEL effort has been submitted to our DOE
Contracting Officer for approval.  We also submitted a CRADA to DOE for approval.  This CRADA involves a
collaboration with STI, Inc to study the design and testing a compact permanent magnet version of our
recirculation bend magnets.

WBS 3 (Beam Physics):
Beam physics activities and opportunities were reported at the CASA staff retreat.  Compaction  management in
FEL drivers was investigated and it was determined that simultaneous bunch length compression during
acceleration and energy compression during energy recovery was possible during recirculation/recovery of
multiple beams, and that for a "magic phase" (value available from your correspondent for a fee) the injected
bunch is imaged at the dump. Provided you're not lasing, of course.

WBS 4 (Injector):
Worked with the cathode shield test setup.  It was found that under vacuum it would still operate but the design
tolerances allowed the mechanism to shift and run stainless against stainless.  We are modifying the design for
better operation. Parts for the cathode stalk were completed by the machine shop and checked.  We will wait to
weld the assembly so we can match machine the stalk seat with the interior cut in the ball when the ball is
completed.  Machining continued on the main gun chamber.

Gun HVPS - The design of the tank and stack mounting continues.

WBS 6 (RF):
Zone 3 - The Cathode Power Supply vendor, Hipotronics, was in this week to replace a current transducer and
to repair a faulty crowbar circuit.  RF power curves were run for position 1 with attenuators up to 6 dB.  This is
to prevent operation on the back side of the power curve.

Zone 4 - Additional RF Control Modules are being calibrated by AES to replace our missing units.

Injector - The second 100kW klystron was tested and approved.  Klystrons S/N - 101 and 102 are to be
shipped from CPI for arrival tonight (Friday).  Unit S/N - 103 is expected to be tested and shipped by the end of
October.  Parts for the HVPS upgrade are still being received.

WBS 8 (Instrumentation):
A new relay box was developed for remote control of equipment by experimenters.  It allows the EPICS
computer to control switches with a given time delay for more accurate timing when more than one function is
being performed simultaneously.

The drawing index on the Laser website has been updated to include many new and updated designs.

Material and test equipment control data bases are being developed to enable us to easily tell our users what we
have available.

Lots of user support this week both for Michelle and the NRL and for the nano-tube experiments.

WBS 9 (Transport):
Optical Chicane Dipoles (GW)
o  The contract was placed with Magnet Enterprises International for 10 of the GW dipoles (6 are required for
    the UV line).  We are preparing the shipment of copper and steel to get them started fast.
Injector Dipoles (DU/DV)
o  The final round of adjustments to the drawings is almost complete.
Arc Dipoles (GY, GX, GQ)
o  Design continued on backchecking the detail drawings of the bend and Reverse Bend Dipoles (GX & GQ)
    with the comments  and corrections.
o  The rework of all drawings of the 180° Dipole (DY) are complete for now subject to last minute adjustments.
Magnet Measurement
o  Work started on the dipole and quadrupole measurement stands
3 inch quad (QX)
o  The assembly quotations are in evaluation.
o  Work continues on cores and coils.
Trim Quad (QT)
o  This magnet is in the procurement cycle.
Sextupole (SF)
o  Work continued on the magnetic model as well as the drawings.
Octupole (OT)
o  DULY Research generated their report on their successful design.  They are now in discussions on how to
    convert the magnetic model into a physical device as an extension of their task order.
Corrector Dipoles (DB, DJ)
o  Milhous Co. is working on the remainder of the coils.
o  The mechanism necessary to mount the DB/DJ coils on to the new large diameter BPMs was signed and sent
    for bid.
Beam Line and Vacuum
o  The exit and entrance of the Optical Chicane chambers were settled this week.
o  The girders and pedestals necessary for the return leg and the hardware for the entire machine entered into a
    procurement cycle.
o  Design work continues on the regions between the cryomodules.
o  Work continued on the injector line stands and vacuum chambers.

WBS 10 (Wiggler):
Initial wiggler measurements indicate that the DC field has been overcompensated.  When a 5.3 G DC field is
added to the measured field the attached trajectory is found.  The trajectory is straight to better than 50 microns
with a peak to peak wiggler amplitude of 1.3 mm.  As expected for an electromagnetic wiggler, the field quality is
exceptional.  The end correction also appears to be quite good with no end pole kick in evidence. Finally, the
trajectory offset from the initial position is less than one wiggler amplitude, which is what was expected.  The
vertical field indicates a probe angle of 1.2 degrees.   We will attempt to fix this and the DC fields before taking
final scans.  So far, no problems with the wiggler field have been seen.  We lost a week of effort on the wiggler
vacuum chamber last week while the vacuum cleaning facility was moved to the test lab addition.  This is delaying
the completion of the vacuum chamber as well.  The final machining of the first chamber is done and the second
chamber is ready for welding.  The first chamber should be complete, including QC and cleaning, next week.
The second chamber should follow by a few days.

               Wiggler Trajectory

WBS 11 (Optics):
We are continuing to order components for the optical cavity.  We've also begun to detail the electrical and water
connections.  We are now turning our attention to other optical components, holding meetings on, the insertable
mirror that diverts the output of the FEL to power and energy detectors.

The first article of our new, near-Brewster window assemblies passed inspection, and is being cleaned for
vacuum service.  The other window assemblies were received this week.  We received new IR cameras from
Raytheon; they are a much-improved version of the cameras we use now to monitor beam position in the optical
transport (at one point in the accelerator vault, in the optics control room, and at the end-of-line.

This week we have been running experiments for the JTO-funded materials studies, and at this writing 15:00
9/28) have irradiated 300 targets, 156 since last Friday.  As we did last week, we supported two visiting
scientists from NRL, who are using their diagnostics to observe the targets during irradiation.  We also hosted
Dr. Dennis Burianek from Lincoln Labs, and irradiated targets of interest to his organization.